The line segment joining A (0, 0) and B (6, -8) is a diameter. Find the equation of a circle.
Answer by Boreal(15235)

The midpoint of the line is the center or (3, -4).
The radius is the distance of each point from the center.
That is the sqrt(3^2+(-4)^2)=sqrt (25), or 5
the equation is (x-h)^2+(y-k)^2=r^2
or (x-3)^2+(y+4)^2=25
Can check to see if each of the two endpoints of the diameter satisfies the equation.